A ball of mass 50 kg is moving with a speed of 2m/s strikes a plane surface at an angle of incidence 45 degrees. the ball is reflected by the plane at an equal angle of reflection at the same speed. calculate (a) the magnitude of the change of momentum of the ball . (b)the change in the magnitude of the momentum of the ball

Grade:11

2 Answers

vikas askiitian expert
13 years ago

let ball is moving in x-y plane & approaching a wall at an angle45 degree which is placed in y-z plane...

initial momentam of ball is Pi =mvcos45 i  + mvsin45(-j)

finally it returnes at same angle & with same velocity so its final momntam is

 Pf = mvcos45(-i) + mvsin45(-j)

change in momentam= Pi - Pi =2mvcos45

                                           =21/2 mv

                                           =100.21/2  or 100sqrt2  ans

approve if u like

Aman Bansal
13 years ago

dear pooja singh ,

initially the ball s trikes theplanem akinganan gleof 45 degrees withthe plane and after the collision the ball makes an angleof45 degrees 

then we canconserve momentum  in both the directions:

x-direction

(a)  magnitude of change in momentum =  50* 21/2    

(b) but change in  magnitude =0

similarly,  

  

y-direction

(a)  magnitude of change in momentum =  0    

(b) but change in  magnitude =0

 

BEST OF LUCK

We are all IITians and here to help you in your IIT JEE preparation.
Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.
Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar..  

AMAN BANSAL